OpenStudy (anonymous):

Which of the following points is not on the graph of y=3{1/3}^x A. (-2,27) B.{3,1/9} C.(-1,8) D.{2,1/3}

OpenStudy (welshfella):

substitute for x in each of the options to see which one fits

OpenStudy (welshfella):

for example , first optionn x = -2 and y = 27 y = 27 3(1/3)^-2 = 3 * 1 / (1/3)^2 = 3 * 1 / (1/9) = ?

OpenStudy (anonymous):

i tried that but it didn't give any exact anwser